### v4

# A Survey of Threats Against Voice Authentication and Anti-Spoofing Systems

Abstract
Voice authentication has undergone significant changes from traditional systems that relied on handcrafted acoustic features to deep learning models that can extract robust speaker embeddings. This advancement has expanded its applications across finance, smart devices, law enforcement, and beyond. However, as adoption has grown, so have the threats. This survey presents a comprehensive review of the modern threat landscape targeting Voice Authentication Systems (VAS) and Anti-Spoofing Countermeasures (CMs), including data poisoning, adversarial, deepfake, and adversarial spoofing attacks. We chronologically trace the development of voice authentication and examine how vulnerabilities have evolved in tandem with technological advancements. For each category of attack, we summarize methodologies, highlight commonly used datasets, compare performance and limitations, and organize existing literature using widely accepted taxonomies. By highlighting emerging risks and open challenges, this survey aims to support the development of more secure and resilient voice authentication systems.
Index Terms:
Privacy, Authentication, Anti-spoofing, Poisoning, Adversarial, Deepfake, Spoofing
I Introduction
Voice authentication is a biometric verification method that leverages the unique characteristics of an individual’s voice to authenticate their identity. It originated as a simple speaker recognition technique in the mid-20th century, relying primarily on handcrafted acoustic features and statistical models such as Gaussian Mixture Models (GMMs). It was initially adopted in restricted environments like forensic analysis and limited-access security systems, where its development was constrained by speech condition variation and limited computational capabilities. However, the landscape changed dramatically with the emergence of deep learning. Over the past decade, deep learning has fundamentally transformed voice authentication, enabling more accurate, robust, and scalable solutions. Advanced models, such as Deep Neural Networks (DNNs), Convolutional Neural Networks (CNNs), and Transformer-based architectures, now extract complex and invariant voice representations, often referred to as speaker embeddings, which significantly improve performance in real-world environments  [1, 2, 3, 4].
Voice authentication plays a critical role in modern security ecosystems due to its non-intrusive nature, ease of implementation, and cost-effectiveness. Its applications span a wide range of domains, from mobile device unlocking to voice assistants, secure banking transactions, smart home systems, and access control in law enforcement and judicial processes [5, 6, 7]. As a contactless solution requiring only a microphone, voice authentication enhances user experience while mitigating risks associated with password leakage, social engineering, and physical spoofing, making it an increasingly important component of multi-factor authentication frameworks [8].
The first documented attack on a voice authentication system was recorded in the 1990s when researchers discovered that replaying a pre-recorded voice sample could effectively bypass early speaker verification systems [9]. These systems lacked robust liveness detection and could not distinguish between a real speaker and a recorded signal, exposing a fundamental weakness that undermined their security in practical scenarios. This initial form of the replay attack became a foundational example in the study of biometric spoofing and highlighted the necessity of building systems capable of discerning not just who is speaking, but also how and when the speech occurs [10, 11].
Despite major advances in deep learning, the sophistication and frequency of attacks against voice authentication systems have significantly increased. As these systems become more widely adopted and technically advanced, various kinds of attacks against these systems also emerge at the same time. One fascinating example occurred when criminals successfully cloned the voice of a company director to defraud a bank in the United Arab Emirates [12, 13]. Leveraging AI voice synthesis, the attackers convincingly impersonated the director through a phone call and reinforced the deception through carefully crafted emails. This coordinated attack ultimately led to the unauthorized transfer of \$35 million. The incident illustrates that, while deep learning has improved voice authentication accuracy, it has also enabled new forms of high-impact, cross-border fraud that traditional security mechanisms struggle to detect and prevent with conventional security controls.
This incident highlights a growing concern: as voice authentication systems become more sophisticated, so do the methods used to compromise them. One critical threat is data poisoning, where adversaries inject malicious or mislabeled data into training datasets or pipelines to compromise model behavior or insert hidden backdoors [14, 15, 16, 17, 18]. While extensively studied in Computer Vision and NLP, poisoning remains underexplored in the audio domain, lacking standardized benchmarks and threat models. Another challenge comes from adversarial attacks, which involve subtle audio perturbations that trick models into misclassification or unauthorized authentication. Though some defenses exist, they often fail under realistic deployment conditions, such as black-box or over-the-air scenarios [19, 20, 21, 22].
Meanwhile, deepfake voice attacks are evolving very quickly. High-fidelity synthetic voices, generated from minimal samples, are now capable of effectively mimicking real users, making it easier to bypass voice authentication systems [23, 24]. Existing countermeasures, including replay and spoofing detectors, are increasingly outpaced by these generative techniques. Even the defensive modules themselves have become targets. Attackers now design adversarial spoofing attacks to evade VAS and spoofing detectors, undermining the reliability of traditional protection measures. This has led to a security arms race in which new defenses trigger more advanced attack strategies, emphasizing the need for holistic, adaptive, and robust approaches to secure voice authentication.
In this review, we provide a comprehensive and systematic survey of the threats targeting voice authentication systems. These threat channels exploit vulnerabilities across the machine learning pipeline—from model training to input inference and security filtering—and challenge the reliability of speaker verification technology. By synthesizing findings from recent literature, we present a structured overview of how these threats operate, how current systems respond, and where critical security gaps remain. This perspective aims to equip researchers and practitioners with a deeper understanding of the evolving adversarial landscape in voice-based authentication.
While prior surveys have explored specific attack classes—such as backdoor attacks [18], adversarial attacks [21], and audio deepfake detection [25, 26, 27]—they have overlooked key areas like broader data poisoning techniques, deepfake attacks themselves, and adversarial spoofing attacks. Moreover, there is currently no unified review that addresses the full spectrum of attacks facing voice authentication systems. Existing works generally treat these threats in isolation, overlooking the interplay and compounding risks that arise when multiple attack strategies are combined. This review fills that gap by providing the first comprehensive study of data poisoning, adversarial, and deepfake synthesis attacks, as well as adversarial spoofing attacks. By integrating these perspectives, we offer an integrated view of the adversarial landscape and its implications for voice-based security systems.
This paper is organized as follows. Section II presents the motivation behind studying threats to voice authentication systems, emphasizing the growing reliance on voice-based security. Section III provides background on voice authentication systems and the typical threat models they face. Section IV examines data poisoning attacks, discussing techniques that compromise training data to subvert model behavior. Section V delves into adversarial attacks, where carefully crafted audio inputs are designed to mislead voice authentication systems during inference. Section VI investigates deepfake attacks, focusing on the creation of synthetic voices and their potential to impersonate legitimate users. Section VII addresses adversarial spoofing attacks, which target anti-spoofing countermeasures to allow malicious audio to bypass security layers. Section VIII surveys prominent datasets, benchmarks, and competitions that support research and evaluation in this domain. Section IX offers a comparison of the different attack strategies discussed. Section X outlines directions for future work, suggesting ways to enhance system robustness. Finally, Section XI provides a conclusion, summarizing the key insights and implications of this review.
II Motivation
As voice authentication becomes popular in financial services, call centers, and personal digital assistants, its appeal as a biometric modality lies in its convenience and natural interaction. However, the growing adoption of voice-based identity systems has drawn the attention of adversaries who exploit both human trust and technical vulnerabilities. In recent years, a ripple has been detected in real-world attacks that challenge the security promises of voice authentication.
One of the most concerning trends is the rise of deepfake voice scams, which leverage AI-generated audio to impersonate individuals and obtain unauthorized access. In 2019, fraudsters cloned the voice of a German CEO to deceive a UK energy executive into sending €220,000 to a fake supplier [28]. In a more sophisticated heist in 2020, attackers combined AI voice synthesis and email spoofing to steal about \$35 million from a UAE bank [29]. These incidents underscore the reality that cloned voices can convincingly pass for legitimate users, especially when combined with social engineering.
Even more recent attacks show the evolving threat landscape. In 2024, a finance employee at Arup Engineering was tricked into transferring \$25 million during a live video call with AI-generated fake executives [30]. Meanwhile, a LastPass employee narrowly avoided a breach after receiving deepfaked audio impersonating the company CEO [31]. Additional demonstrations have further shown how adversarial audio samples and even analog strategies, such as speaking through a PVC pipe, can subvert voice recognition systems or bypass liveness detection [32, 33].
Collectively, these real-world cases reveal that the threat is no longer theoretical. From synthetic voices to adversarial bypasses, attackers have started exploiting the very attributes that once made voice authentication seem secure. As a result, a comprehensive survey is urgently needed to evaluate these vulnerabilities and guide future research toward more robust and resilient voice authentication systems.
III Background
III-A List of Acronyms
To facilitate understanding, this subsection provides a list of acronyms commonly used throughout the document (see Table I).
TABLE I: List of Acronyms
Acronym Definition
SV Speaker Verification
ASV Automatic Speaker Verification
SRS Speaker Recognition System
SRE Speaker Recognition Evaluations
CSI Closed-set Identification
OSI Open-set Identification
VAS Voice Authentication Systems
VC Voice Conversion
TTS Text-to-Speech
ADD Audio Deepfake Detection
DFDC Deepfake Detection Challenge
PAD Presentation Attack Detection
PF Partial Fake
CMs Anti-spoofing countermeasures
GMMs Gaussian Mixture Models
UBM Universal Background Model
DNN Deep Neural Network
CNN Convolutional Neural Networks
k-NN k-Nearest Neighbors
GAN Generative Adversarial Network
MFCC Mel-Frequency Cepstral Coefficients
GFCC Gammatone Frequency Cepstral Coefficients
SNR Signal-to-Noise Ratio
RIR Room Impulse Response
FIR Finite Impulse Response
DET Detection Error Tradeoff
EER Equal Error Rate
ASR Attack Success Rate
FGSM Fast Gradient Sign Method
BIM Basic Iterative Method
PGD Projected Gradient Decent
NES Natural Evolution Strategy
UAPs Universal Adversarial Perturbations
III-B Voice Authentication Systems Overview
Voice Authentication Systems (VAS) have evolved significantly over the years, transitioning from traditional statistical models to advanced deep learning-based architectures. Understanding the architecture and taxonomy of these systems is essential for identifying their vulnerabilities, operational contexts, and security challenges. This section provides a detailed overview of the underlying architectures and the different types of voice authentication tasks.
III-B1 System Architectures
A typical speaker recognition system (SRS) operates in three main stages, as illustrated in Figure 1:
  1. 1.
    Training phase: The system is trained on a large dataset of voices to extract speakers’ discriminative features and learn a robust embedding space, often using deep learning models.
  2. 2.
    Enrollment phase: Each user’s voiceprint is recorded through multiple samples and transformed into a reference embedding, which is securely stored.
  3. 3.
    Recognition phase: During authentication, a new voice sample is captured, embedded, and compared to the enrolled reference using similarity scoring and thresholding strategies.
Figure 1: Speaker Recognition System Overview: (a) Overview in the past. (b) Current system – showing the modernized process where models have evolved with advancements in AI and data handling.
In traditional systems (Figure 1.a), the process begins with pre-processing and feature extraction, followed by training a Universal Background Model (UBM)[34]. Speaker-specific models are derived from the UBM during enrollment, and verification is performed using statistical techniques such as Gaussian Mixture Models (GMMs)[35]. Models like i-vector [36] significantly improved system compactness and efficiency by encoding speaker characteristics into low-dimensional embeddings. However, this multi-stage architecture is sensitive to environmental noise and variability in speaker conditions.
Modern systems (Figure 1.b) have adopted end-to-end deep learning architectures. These systems unify pre-processing, feature extraction, and speaker modeling into a single pipeline. Rather than constructing explicit speaker models, embeddings are extracted directly from speech signals and compared using similarity metrics like cosine similarity or learned distance functions [37, 38]. While these models offer improved accuracy and scalability, they still face challenges such as environmental noise, microphone variability, and resistance to spoofing attacks (e.g., replay, synthesis, and voice conversion).
III-B2 Voice Authentication Tasks
Voice authentication tasks can be categorized along three primary dimensions: the recognition task type, the interaction mode, and the content dependency.
  1. 1.
    Recognition Task Types:
    • Speaker Verification (SV): A one-to-one task that verifies the claimed identity of a speaker by comparing the input voice sample to a stored reference embedding. Authentication is accepted if the similarity score exceeds a defined threshold.
    • Closed-set Identification (CSI): A one-to-many classification task that identifies the speaker from a known set of enrolled users. It does not support rejection; the system must select one of the enrolled identities.
    • Open-set Identification (OSI): Similar to CSI but includes a rejection option, allowing the system to classify a speaker as “unknown” if the similarity score does not meet the threshold.
    • Speaker Diarization: Focuses on the question “Who spoke when?” by segmenting and clustering an audio stream according to different speakers. This task is crucial for applications like meeting transcription or call analytics.
  2. 2.
    Interaction Modes:
    • Active Systems: Require users to consciously participate by speaking prompted phrases or performing voice-based actions. These systems benefit from controlled conditions, improving accuracy.
    • Passive Systems: Operate in the background during natural conversations, offering unobtrusive authentication. However, they face greater challenges related to noise, variability, and spoofing.
  3. 3.
    Content Dependency:
    • Text-dependent Systems: Rely on predefined phrases spoken by the user. Consistent wording aids system accuracy but limits user flexibility.
    • Text-independent Systems: Allow for authentication using any speech content, offering a more natural experience and enabling passive use cases.
III-C Attacks Definitions in Voice Authentication
Despite significant advances, voice authentication systems remain vulnerable to a diverse range of attacks that target both the training and inference phases. These attacks are often motivated by the goal of impersonating users, bypassing authentication, or degrading system performance.
To better understand the evolving focus of the research community on these threats, Figure 2 shows the cumulative number of research studies published over time that specifically address attacks against voice authentication systems.
Figure 2: Cumulative number of research studies on attacks against voice authentication systems.
Data poisoning attacks involve injecting carefully crafted malicious samples into the training set to manipulate the model’s behavior. A notable subclass is backdoor attacks, in which an attacker introduces a specific pattern or trigger during training that subsequently can be activated at inference time. These attacks are particularly concerning in federated learning, where malicious participants can corrupt the global model by compromising the client model. Defending against poisoning requires robust training techniques such as outlier detection, differential privacy, or adversarial training.
Adversarial attacks exploit the vulnerability of deep neural networks by adding imperceptible perturbations to audio inputs that cause misclassification. Several techniques have emerged, including additive noise, psychoacoustically-aware perturbations that remain imperceptible to humans, and universal adversarial perturbations (UAPs) that generalize across inputs.
Figure 3: Comprehensive taxonomy of attacks on voice authentication systems, categorized by key factors including attacker knowledge, objectives, access level, perturbation type, transferability, medium, delivery channel, and attack vector.
Deepfake voice attacks involve the use of speech synthesis (TTS or VC) to mimic the vocal identity of a target speaker. Modern deepfake tools can operate in real time, increasing the feasibility of live impersonation. The success of these attacks depends on the quality of training data and synthesis models. Recent advancements in zero-shot voice cloning enable attackers to generate convincing deepfakes with minimal reference audio, which lowers the barrier to entry. Currently, researchers produce anti-spoofing countermeasures (CMs) to prevent such attacks [27].
After designing anti-spoofing models to prevent deepfake audio attacks, attackers have adapted by launching adversarial spoofing attacks against both the VAS and the CMs. These attacks typically involve adding subtle perturbations to AI-generated audio, crafted by speech synthesis models, to fool the authentication system and spoofing detectors.
III-D Attack Taxonomies
The landscape of attacks in voice authentication can be systematically categorized along several dimensions, as illustrated in Fig. 3.
Attacks are often characterized by the attacker knowledge they have access to:
  • White-box: The attacker has complete access to model architecture, training data, and parameters.
  • Gray-box: The attacker has partial knowledge, such as access to speaker embeddings, loss functions, or system design.
  • Black-box: The attacker has no internal system knowledge and relies solely on input-output behavior, often using query-based attacks or surrogate models.
Attack objectives in VAS depend on the attack goals pursued by the adversary:
  • Targeted Attacks: The attacker aims to impersonate a specific legitimate user. This includes impersonation attacks such as replay, voice conversion, or deepfake-based synthesis, where the goal is to bypass authentication by pretending to be a target speaker.
  • Untargeted Attacks: The attacker’s goal is to cause system failures, like incorrect classification, such as misclassification, rejection of legitimate users, or degradation of model confidence.
The type of attack is influenced by the access level:
  • Training phase: Enables attacks during the model training phase, including data poisoning, backdoors, or embedding adversarial triggers into training data.
  • Inference phase: Involves attacks during model deployment or usage, such as adversarial perturbations, spoofing, and query-based attacks that exploit model behavior.
The generality of an attack is described by its attack universality:
  • Universal Attacks: One crafted perturbation or method can fool multiple inputs or models.
  • Customized or Targeted Attacks: These attacks are tailored to a particular model or input. They often require detailed knowledge of the model architecture or parameters and typically achieve high success rates against that specific target; however, they may not generalize well to other models or inputs.
  • Transferable Attacks: Attacks crafted on one surrogate model work against another target model, which is called cross-model vulnerability.
Transferable attacks exploit adversarial examples that remain effective across different settings:
  • Cross-Model: Attacks generated on one model architecture succeed on others with different designs.
  • Cross-Dataset: Perturbations crafted on one dataset remain effective on another, despite differences in data distribution.
  • Cross-Protocol: Attacks transfer between tasks such as verification and identification.
  • Cross-Environment: Adversarial audio retains its impact under noise, channel shifts, or hardware variation.
  • Cross-Speaker: Attacks spoof multiple user profiles beyond the original target.
  • Cross-Domain: Transferability occurs between synthetic and natural speech inputs.
The delivery method of adversarial input is called the attack delivery channel:
  • Over-the-line: Inputs are delivered digitally through online channels such as file uploads, VoIP calls, APIs, or compromised communication links.
  • Over-the-air: Adversarial audio is played aloud in a physical space with the intent that a microphone will capture it. This includes traditional replay attacks, adversarial TTS played through speakers, and inaudible signals such as ultrasonic commands.
The method of training data compromise falls under attacker vectors in poisoning taxonomy:
  • Dirty-label Poisoning: The attacker manipulates both the input $(x_{i})$ and the corresponding label $(y_{i})$, often injecting incorrect labels to misguide learning.
  • Clean-label Poisoning: Only the input $(x_{i})$ is subtly altered, while the label remains correct. These attacks are stealthier and harder to detect since the data appears valid.
The form of adversarial input is referred to as the attack medium:
  • Replay Attacks: Use of pre-recorded audio from a target speaker, typically played back via a speaker or other device to simulate a live voice.
  • Synthetic Speech: Generation of voice samples using AI-based techniques (TTS or VC) to imitate a target speaker’s voice.
  • Adversarial Perturbations: crafted modifications to benign audio signals that cause misclassification without being noticeable to humans.
  • Environmental Manipulation: Attacks involving the physical space, such as injecting perturbations through room acoustics, exploiting microphone characteristics, or introducing background signals to interfere with the system.
III-E Threat Models in Voice Authentication Attacks
Voice authentication systems are vulnerable to a variety of sophisticated threats, each exploiting different stages of the machine learning lifecycle or system architecture. Figure 4 illustrates the relationship between these attack types.
Figure 4: Attacks Threat Models (a) Data Poisoning Attacks. (b) Adversarial Attacks. (c) Deepfake Audio Attacks. (d) Adversarial Spoofing Attacks.
III-E1 Data Poisoning Threat Models
Data poisoning attacks target the training phase of VAS by injecting malicious or manipulated data into the model’s training pipeline to corrupt its learned representations or behavior.
As discussed in Section III-D, the attacker’s knowledge can vary: in a white-box setting, they may craft highly optimized poisoned inputs using full model access; in a gray-box scenario, poisoning is guided by partial system knowledge; and in a black-box setting, attacks rely on limited feedback or surrogate datasets. Poisoning goals may be targeted, such as embedding a backdoor to impersonate a specific user, or untargeted, aiming to degrade the model’s overall performance. Attack vectors include both clean-label and dirty-label strategies, depending on whether the injected samples are correctly or incorrectly labeled.
Further technical details and relevant literature will be discussed in Section IV.
III-E2 Adversarial Threat Models
Adversarial attacks exploit vulnerabilities during the inference phase by introducing carefully crafted perturbations to input audio. These perturbations are typically imperceptible to humans and constrained under $L_{p}$-norm bounds (e.g., $_{2}, LL_{\infty}$), allowing them to bypass detection while manipulating model predictions.
According to the taxonomy in Section III-D, adversarial attacks may occur under white-box conditions, where the attacker leverages full model access to compute gradients; gray-box settings, with partial system knowledge; or black-box scenarios, relying on queries or transferability. The objective may be targeted, forcing the system to accept an adversary as a specific user, or untargeted, simply inducing misclassification or denial of access. Effective adversarial examples maintain a trade-off between attack success and perceptual naturalness.
Further technical details and relevant literature will be discussed in Section V.
III-E3 Deepfake Voice Threat Models
Deepfake voice attacks exploit generative models to synthesize speech that mimics a target speaker, aiming to deceive VAS. The success of these attacks depends largely on the attacker’s available resources.
The attacker’s goal is typically targeted, seeking to impersonate a specific speaker by generating audio that meets or exceeds the similarity threshold $\tau$ required for authentication. Common techniques include VC, which transforms one speaker’s voice to sound like another’s, and TTS, where new phrases are generated using a cloned voice profile. These attacks challenge a system’s ability to distinguish between genuine and synthetic input, particularly as deep generative models continue to improve.
Further technical details and relevant literature will be discussed in Section VI.
III-E4 Adversarial Spoofing Threat Models
Some of the most critical threats emerge when adversarial perturbations are designed to deceive both VAS and anti-spoofing CMs. These attacks typically utilize the output of TTS or VC systems, which are already capable of targeting VAS, and further adapt them to bypass anti-spoofing CMs as well.
These attacks can follow a white-box paradigm, where the adversary has access to gradients of both the speaker recognition model $f_{\theta}$ and the spoof detector $d_{\phi}$, allowing for coordinated optimization across both targets. These dual-purpose attacks typically involve minimizing a joint loss function: $\mathcal{L}_{\text{auth}}+\lambda\mathcal{L}_{\text{spoof}}$, where $\lambda$ adjusts the emphasis between breaking authentication and bypassing spoof detection. In black-box or gray-box settings, attackers use surrogate models or query-based methods to approximate behavior. The attack goal may be either targeted or untargeted, depending on whether the adversary seeks impersonation or generalized evasion. This multi-objective threat landscape underscores the need for robust, multi-layered defense strategies.
Further technical details and relevant literature will be discussed in Section VII.
IV Data Poisoning Attacks
TABLE II: Summary of Data Poisoning and Backdoor Attacks against Voice Authentication Systems
Study Year Mechanism Target Poisoning Rate ASR Notes / Limitations
Untargeted Data Poisoning
[39] 2021 Adv. Examples CNN-SV Requires pre-trained model and full dataset access
[40] 2021 Unlearnable Examples Audio CNNs Vulnerable to preprocessing/augmentation
[41] 2022 Clean-label flip MFCC-based models $\geq 10\%$ Less effective on ResNet/FFT, Kaldi shows robustness
[42] 2025 Synthetic Pop Noise GFCC + SVM 20% 95% White-box, pop-noise specific, mitigated by suppression
Targeted Data Poisoning
[43] 2018 Feature Collision MFCC-based models Classical attack, less explored in audio domain
[44] 2023 Enrollment Injection DeepSpeaker 10% Train. - 50% Enroll 90% Unrealistic assumptions about enrollment data access
[45] 2024 Low-rate Label Flip LibriSpeech, VoxCeleb 5% Assumes attacker controls training, limited in deployed systems
Backdoor Attacks
[46] 2018 Neuron Trigger Neural Networks 95.5–100% Full white-box access required
[47] 2021 Blending, Parameter Speaker Recognition Models 98.66% Detectable triggers, poor transfer to text/tabular data
[48] 2021 Clustering Trigger SV Systems 15% 63.5% Limited stealth at high rates, cluster sensitivity
[49] 2022 Ultrasonic Frequency Trigger Frequency-sensitive Models ¡5% 95–100% Hardware sensitivity, fails under downsampling
[50] 2023 Voice Conversion Trigger VoxCeleb1-based Models 1% 99.94% Fails for mismatched voices, task-dependent performance
[51] 2024 Emotional Prosody Trigger ResNet, X-vector, ECAPA 5–10% 98.9% Emotion-dependent, pruning reduces efficacy
[52] 2025 Backdoor Defense Spectrogram-based Detectors N/A Reveals PBSM stealth limits, preprocessing counters attack
Data poisoning attacks are a critical security concern for VAS, where adversaries manipulate training data to compromise a model’s integrity or availability [15, 16, 53]. With the increasing reliance on third-party data sources, including cloud-based services and web scraping, the risk of unintentionally incorporating poisoned data into training pipelines has increased significantly, particularly in the audio domain, where model training is computationally expensive and data-hungry [53, 15]. A summary of these attacks is illustrated in Table. II.
IV-A Untargeted Data Poisoning Attacks
Untargeted data poisoning attacks aim to degrade a model’s overall performance without targeting specific outputs. Fowl et al. [39] demonstrated this by injecting adversarial examples into training sets. Their method uses crafted Projected Gradient Descent (PGD) perturbations on a substitute CNN to generate semantically valid but misleading inputs. Although this was effective in lab settings, it requires access to the pretrained model and full-batch training data, which often breaks down in streaming or privacy-preserving voice data pipelines, limiting real-world applicability.
Huang et al. [40] proposed an alternative approach using unlearnable examples, which suppress informative features in training data, disrupting representation learning in audio models. Although this clean-label method cleverly avoids detection through label inspection, its effectiveness heavily depends on the model’s inability to recover signal through preprocessing or data augmentation, standard practices in modern voice pipelines. This reliance makes the attack less transferable.
WaveFuzz by Ge et al. [41] takes clean-label poisoning a step further by changing Mel-Frequency Cepstral Coefficients (MFCC) features to mislead speech models without affecting human perception. Their method lowers accuracy in end-to-end training, but it relies on certain feature extraction techniques, like MFCC. It also has less effect on stronger architectures, such as ResNet-18 and Kaldi [54], which shows its limited ability to generalize. Additionally, the need to poison at least 10% of the training data brings up concerns about practicality in real-world use.
A more targeted attack is SyntheticPop [42], which exploits liveness detection systems that use pop noise, like VoicePop [55], by embedding low-frequency sine-wave “pops” into training samples. This manipulation at the feature level causes Gammatone Frequency Cepstral Coefficients (GFCC)-based detectors to fail. The accuracy drops from 69% to 14% with only 20% poisoning. While this approach is impressive, it assumes white-box access to the target’s audio pipeline. It does not work against systems that use different feature sets or basic denoising, which lowers its threat level in more diverse or well-defended environments.
Across these works, a pattern emerges. Untargeted poisoning methods for audio depend heavily on certain model assumptions, such as known architectures, fixed features, and large-scale poisoning. This reliance weakens their transferability and realism. The field needs more flexible and discreet poisoning strategies that consider audio-specific issues like preprocessing pipelines, limited labels, and data streaming.
IV-B Targeted Data Poisoning Attacks
Targeted data poisoning attacks aim to manipulate machine learning models to misclassify specific targets at inference time, without degrading overall performance. These attacks pose a critical risk to biometric systems, where identity spoofing can have serious consequences.
One influential method is the feature collision poisoning introduced by Shafahi et al. [43]. In this approach, adversarial examples are created to resemble target samples in the feature space while still looking harmless in the input space. This technique achieved a 100% success rate in transfer learning scenarios using a single poisoned sample, with a negligible accuracy loss. However, its scalability to end-to-end models is limited; achieving comparable success required up to 50 poison instances, enhanced using a 30% opacity watermark of the target, highlighting a trade-off between attack stealth and practicality.
Expanding into the voice domain, Li et al. [44] showed a two-phase attack on speaker recognition models like DeepSpeaker [38]. They used label flipping during training and later mixed attacker samples with victim identities in the enrollment process. This method achieved up to 90% success while keeping overall accuracy high. However, their threat model assumes access to the training pipeline and uses a large number of enrollment samples (10). This does not reflect real-world practices, where systems usually use fewer samples (1–3) [56, 57]. Additionally, the attack’s effectiveness depends on ideal conditions, which may not be present in production systems. This raises doubts about its practical use.
Addressing these limitations, Mohammadi et al. [45] proposed a more practical evaluation of poisoning threats. They focused on stealthy, small-scale poisoning with 5% data contamination. They introduced a hybrid CNN and k-Nearest Neighbors (k-NN) defense that exceeded earlier defenses (The Guardian [44]) in detection accuracy on real-world datasets, such as VoxCeleb [58]. Their approach balances robustness and usability, but it still assumes that the attacker can inject poisoned samples labeled as the victim. This assumption may not match typical production systems where training happens offline and is separated from user-specific data.
In summary, targeted poisoning attacks have produced strong results in controlled settings. However, there are major gaps between these situations and real-world systems. The success of these attacks mainly relies on access to the training process and the inner workings of the model. This highlights the need to create threat models that consider practical deployment environments.
IV-C Backdoor Attacks
Backdoor attacks in machine learning involve inserting harmful behaviors into a model during training. These behaviors activate only when specific, often subtle, triggers are present. Unlike data poisoning, backdoor attacks aim to maintain high accuracy on clean inputs while secretly allowing targeted misclassification. This threat is especially pronounced in speaker verification and identification systems, where backdoors can silently override identity checks.
IV-C1 Trigger Mechanisms
Early work by Liu et al. [46] introduced a foundational Trojaning method, embedding triggers by manipulating high-activation neurons and reverse-engineering synthetic training data. This attack achieved near-perfect success across domains (including face recognition and driving systems) with minimal computational overhead and no loss in model accuracy. However, it relied on full white-box access, which limits applicability in production environments.
Building on this, Xu et al. [47] proposed a broader taxonomy of backdoor strategies—modification, blending, and parameter/latent attacks—each targeting different modalities. While successful in vision and speech domains, attacks like blending were less effective due to feature interference, and parameter-based methods struggled with non-vision tasks. Crucially, these attacks required access to the training dataset, reducing their stealth in practical deployment settings.
IV-C2 Audio-Specific Triggers
As attacks advanced, researchers explored novel, audio-specific triggers that exploit unique acoustic properties. For instance, Zhai et al. [48] designed a clustering-based approach for speaker verification, using low-energy spectral noise per cluster. Their attack retained stealth (minimal Equal Error Rate (EER) elevation) while outperforming traditional backdoors like BadNets. However, it traded off simplicity for operational complexity, requiring multiple trigger variants and increased inference-time burden.
A more subtle and innovative approach was introduced by Schoof et al. [51] in EmoBack, which utilized emotional prosody (e.g., pitch and speech rate) as a dynamic trigger. This strategy achieved an Attack Success Rate (ASR) up to 98.9% while preserving clean accuracy. Yet, its performance varied by emotion and model architecture, and it required higher poisoning rates (5–10%), reducing stealth. Defensive pruning degraded the attack partially, underscoring the ongoing cat-and-mouse nature of mitigation.
Pushing stealth to the extreme, Zhang et al. [50] leveraged voice conversion using FreeVC [59] to poison training data with speaker-morphed utterances, achieving 99.94% ASR with only 1% poisoning. These samples preserved the original semantics while replacing the speaker identity—a significant step toward imperceptible attacks. However, their generalizability was constrained: effectiveness sharply dropped when voiceprints did not match across train and test speakers, and success remained architecture-dependent (e.g., RawNet3 ¿ X-vectors).
IV-C3 Inaudible Triggers
In parallel, Cai et al. [49] introduced inaudible triggers via ultrasonic frequencies (18–22 kHz) and pitch modifications (PBSM), exploiting psychoacoustic masking. These triggers were effective (ASR up to 100%) yet vulnerable to environmental noise and hardware variability. Follow-up analysis by Mohammadi et al. [52] demonstrated that these spectral attacks can bypass traditional defenses but fail under common preprocessing techniques like downsampling or non-linear transformations.
This growing class of non-semantic triggers—from environmental sounds (e.g., keyboard clicks) to emotional embeddings—offers high success rates but raises concerns about practical deployment. Triggers must survive real-world transformations and remain undetectable across devices, which is not always feasible. Emotional triggers, for example, may be misclassified due to natural variation, leading to false positives.
V Adversarial Attacks
TABLE III: Comparison of Key Adversarial Attacks on Speaker Recognition Systems.
Reference Year Attack Type Target System Method Key Results Key Findings / Limitations
Foundational Work
Hautamäki et al. [60] 2013 Human Mimicry GMM-UBM, i-vector Professional impersonation FAR: 9.03% $\rightarrow$ 11.61% Less effective than computational synthetic methods.
White-Box Attacks
Kreuk et al. [61] 2018 Gradient-based, Transferable End-to-end Neural Nets FGSM on MFCCs Accuracy: 85.5% $\rightarrow$ 37.5%
FPR: 4.88% $\rightarrow$ 94.63%
Demonstrated high transferability across models and datasets.
Li et al. [62] 2020 Gradient-based, Transferable GMM-UBM, i-vector FGSM on feature space FAR $\approx$ 100% (white-box)
EER up to 50% (black-box)
Achieved high success with imperceptible noise; transferable.
Li et al. [63] 2020 Over-the-Air (OTA) x-vector Gradient optimization + RIR estimation ASR: 98% (digital), 50% (OTA) Focused on practical, physically robust perturbations.
Joshi et al. [64] 2021 Iterative Gradient-based x-vector BIM, PGD, CW attacks Accuracy $\rightarrow$ 0% at $\ell_{\infty}\leq 0.001$ Showed extreme vulnerability to minimal perturbations.
Black-Box Attacks
Zhang et al. [65] 2020 Score-based, Stealth i-vector, x-vector Psychoacoustic masking (VMask) FAR: 3.9% $\rightarrow$ 88.2% Perturbations are psychoacoustically masked to be imperceptible.
Chen et al. [22] 2021 Score-based, Targeted (FakeBob) GMM-UBM, i-vector Natural Evolution Strategy + BIM ASR: 90.5% (i-vector) Effective but requires thousands of queries per sample.
Chen et al. [66] 2023 Query-free, Transfer-based (QFA2SR) Commercial APIs (Azure) Surrogate ensemble + Knowledge Distillation ASR up to 89.3% Eliminates query cost but requires costly surrogate model training.
Physical Attacks
Chen et al. [67] 2022 Transfer-based (AS2T) SV, CSI, OSI systems Multi-task learning + RIR simulation ASR up to 98.5% Demonstrates strong generalization across different speaker recognition tasks.
Chen et al. [68] 2022 Real-time (PhyTalker) Real-time SRS Time-domain phoneme injection Targeted ASR: 85% (OTA) Real-time physical attack without requiring model access.
Adversarial attacks introduce subtle, often imperceptible, perturbations into speech signals to mislead Speaker Recognition Systems (SRS). Unlike spoofing attacks, which typically involve synthesized or replayed speech, adversarial attacks directly manipulate the input data to exploit and alter the model’s decision boundaries. As speaker recognition technology has transitioned from traditional methods, such as Gaussian Mixture Model-Universal Background Model (GMM-UBM) pipelines, to deep learning-based architectures like x-vectors [37] and ECAPA-TDNN [69], the sophistication of attack strategies has advanced in parallel, evolving from purely digital manipulations to practical over-the-air implementations. Table III summarizes key adversarial attacks on speaker recognition systems.
V-A Foundational Work
The concept of adversarial audio in speaker verification was first explored by Hautamäki et al. [60], who employed professional impersonators to target Speaker Verification (SV) systems. Although the false acceptance rate (FAR) increased only slightly from 9.03% to 11.61%, these mimicry attacks demonstrated that human impersonation poses a less significant threat than computational, synthetic methods.
A pivotal shift occurred with the work of Kreuk et al. [61], who applied the Fast Gradient Sign Method (FGSM) to generate targeted adversarial perturbations for end-to-end neural models. This attack reduced accuracy from 85.5% to 37.5% on the YOHO dataset [70] and increased the false positive rate from 4.88% to 94.63%. Furthermore, the adversarial examples demonstrated transferability across different datasets and model architectures, highlighting a systemic vulnerability in neural SRS.
V-B White-Box Attacks
Li et al. [62] first extended these gradient-based attacks to traditional systems, including GMM-UBM and i-vector models. By applying FGSM to the feature space, they increased the FAR to nearly 100% using imperceptible perturbations. Building on this, Li et al. [63] designed practical attacks against the state-of-the-art x-vector system, with both untargeted and targeted objectives. A key innovation was the integration of an estimated Room Impulse Response (RIR) to craft perturbations robust enough to withstand real-world playback distortions. The approach achieved a 98% attack success rate (ASR) in digital simulations and 50% in over-the-air physical playback tests, marking a significant advance in physically realizable adversarial attacks.
Further exploring these vulnerabilities, Joshi et al. [64] analyzed modern x-vector-based systems under white-box conditions. They demonstrated that iterative attacks such as the Basic Iterative Method (BIM), Projected Gradient Descent (PGD), and Carlini-Wagner (CW) could reduce classification accuracy to nearly zero, even with a minimal perturbation budget ($\ell_{\infty}\leq 0.001$). In parallel, Zhang et al. [65] developed VMask, which utilizes psychoacoustic masking to craft perturbations that are imperceptible to humans but effective against both commercial and open-source systems. However, its real-world applicability was hindered by sensitivity to playback hardware and environmental noise. These findings collectively underscore the fragility of deep speaker representations against a variety of white-box threats.
V-C Black-Box Attacks
The FakeBob attack [22] demonstrated a practical black-box approach by combining the Natural Evolution Strategy (NES) and BIM to achieve a targeted ASR exceeding 90% against i-vector systems. However, its practicality was limited by its high query cost, requiring thousands of model queries per sample.
To address the high query requirement, Chen et al. [66] proposed QFA2SR, a query-free black-box attack. It leverages ensemble-based surrogate models and knowledge distillation to generate adversarial examples without direct access to the target system. This method achieved an 89.3% ASR against commercial APIs, including Microsoft Azure, but its reliance on complex surrogate model training requires significant computational resources.
V-D Physical Over-the-Air Attacks
To bridge the gap between digital simulations and real-world applications, Chen et al. [67] proposed AS2T, a generalizable transfer-based framework targeting SV, CSI, and OSI. By utilizing multi-task learning and RIR simulation, AS2T achieved an ASR of up to 98.5% in white-box settings and demonstrated strong cross-task generalization.
PhyTalker [68] extended adversarial attacks to the real-time physical domain by injecting precisely timed, phoneme-level perturbations during live playback. PhyTalker achieved an 85% targeted ASR in over-the-air t
VI Deepfake Attacks
TABLE IV: Summary of Key Deepfake Attacks on Voice Authentication Systems.
Reference Year Attack Type Target Method Key Results Limitations / Notes
Saxena et al. [71] 2015 Full Voice Conversion ASV & Human Listeners Voice morphing using a few minutes of source audio ASV bypass rate: 80–90%
Human accuracy $\approx$ 50%
Used early-generation tools; lacked environmental or cross-lingual testing.
Alali et al. [72] 2025 Partial Fake Speech SRS & Human Listeners Zero-shot TTS (YourTTS) and VC (VITS) for partial synthesis High success rates; reduced detection accuracy for both systems and humans. Not evaluated in noisy or multilingual settings; defenses not thoroughly assessed.
Deepfake speech attacks exploit advanced voice synthesis techniques, such as Voice Conversion (VC) and Text-to-Speech (TTS), to impersonate users in Speaker Recognition Systems (SRS). These attacks can bypass both automated authentication systems and human listeners by generating speech that closely mimics a target speaker’s voice, often using minimal training data. Unlike traditional replay spoofing, deepfake attacks can fabricate entirely new utterances, significantly expanding the threat landscape. As generative models become more powerful and accessible, these attacks pose a growing challenge to voice-based security. A summary of key deepfake attacks is presented in Table III.
Scope Note: This review focuses specifically on studies that evaluate speech synthesis technologies as attack vectors against speaker recognition systems. While the field of generative audio synthesis is vast, encompassing foundational models like Tacotron 2 [73] and advanced architectures like VALL-E [74], we exclude works that do not provide an empirical assessment of their models in an adversarial context. This scope ensures our analysis remains centered on demonstrated vulnerabilities in voice biometrics. For a broader overview of modern speech synthesis technologies, see [75, 76, 73, 77, 74, 3, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 59, 96, 97].
VI-A Foundational Work
One of the earliest empirical demonstrations of AI-driven voice spoofing was presented by Saxena et al. [71]. Using a commercially available voice morphing tool, they synthesized a target speaker’s voice from just a few minutes of enrollment data. The resulting audio was evaluated against both automated speaker verification systems and human listeners. The results were striking: state-of-the-art SRS accepted 80–90% of the spoofed attempts, while human listeners correctly identified the fakes only about half the time.
This study served as an early warning of the potency of voice conversion attacks and underscored the limitations of both machine-based and human auditory verification. It also highlighted the urgent need for robust anti-spoofing and liveness detection mechanisms as high-fidelity synthesis technologies became more widespread.
VI-B Partial Fake (PF) Attacks
Recent research has explored the threat of partial fake (PF) speech, where only specific segments of an utterance are synthetically generated while others remain authentic. Alali et al. [72] investigated PF attacks using modern zero-shot TTS and VC models, including VITS and YourTTS. Their findings revealed that both humans and SRS models are highly susceptible to these nuanced manipulations. Because they preserve characteristics of genuine speech, partial fakes are often more challenging to detect than fully synthetic samples, particularly in zero-shot scenarios that require no speaker-specific fine-tuning.
Although the study demonstrates high attack success rates, it also highlights several research gaps. The attacks were not evaluated under noisy or other challenging real-world conditions, such as those involving channel variability or background speech. Consequently, the robustness of these attacks and potential countermeasures in practical deployment scenarios remains an open area for investigation.
VII Adversarial Spoofing Attacks
TABLE V: Summary of Adversarial Spoofing Attacks on Voice Authentication and Anti-Spoofing Systems.
Reference Year Access Method Target Effect on Target Notes / Limitations
White-Box Attacks
Liu et al. [98] 2019 White-box Gradient-based perturbations (FGSM, PGD) PAD Increases error rate to >>85% Requires model gradient access; bounded by small perturbation norm.
Gomez-Alanis et al. [99] 2022 White-box ABTN (Adversarial generator with PAD+ASV loss) PAD, ASV Increases attack EER by 27–51% (absolute) Generalizes better than FGSM while preserving speaker identity.
Gomez-Alanis et al. [100] 2022 White-box GAN-based spoofed speech generation (GANBA) PAD, ASV Reduces ASV accuracy to <<20%; degrades PAD Computationally intensive; requires extensive training.
Black-Box Attacks
Hai et al. [101] 2023 Black-box Query-based perturbation of background/silence PAD Achieves >>80% ASR while preserving speaker identity Attack is query-heavy and tailored to specific PAD vulnerabilities.
Kassis et al. [102] 2023 Black-box Over-the-air signal processing attack ASV, CM Bypasses multiple systems via physical playback Requires >>15 min of target data; less effective on robust CMs.
Panariello et al. [103] 2023 Black-box (Universal) Universal FIR adversarial filter (Malafide) PAD Enables real-time, universal bypass of a target PAD Filter is universal for one PAD but may not generalize to others.
Todisco et al. [104] 2024 Black-box (Universal) Universal nonlinear filter (Malacopula) ASV Enhances voice match of spoofed audio to bypass ASV Targets the ASV system directly, not the PAD; physically viable.
Anti-spoofing countermeasures (CMs), also known as Presentation Attack Detection (PAD), are a critical defense layer in Speaker Verification (SV) pipelines, designed to distinguish bona fide speech from spoofed audio. However, recent studies have demonstrated that these systems are themselves highly vulnerable to adversarial attacks. These attacks introduce crafted, often imperceptible, perturbations to degrade the reliability of CMs, thereby increasing the false acceptance of synthetic or replayed speech. Table V presents a summary of these attacks.
VII-A White-Box Attacks
Early work by Liu et al. [98] applied white-box adversarial techniques, such as the FGSM and PGD, against PAD models. Even minimal, imperceptible perturbations (i.e., small $\epsilon$) increased detector error rates to as high as 85–95%. These results revealed that, with access to model gradients, an attacker can nearly eliminate a PAD system’s ability to detect spoofed speech, effectively reducing its performance to that of random guessing.
Building on this, Gomez-Alanis et al. [99] proposed a trainable adversarial generator named ABTN (Adversarial Biometrics Transformation Network), designed to optimize perturbations using both PAD and Automated Speaker Verification (ASV) losses. ABTN outperformed traditional FGSM by generating perturbations that generalized better across spoof types while preserving the speaker’s identity. This allowed the attack to fool the PAD system without compromising the ASV system’s ability to verify the speaker. The method improved the attack Equal Error Rate (EER) by an absolute 27–51% on the ASVspoof 2019 logical and physical access protocols [105].
A different approach, GANBA [100], introduced a conditional Generative Adversarial Network (GAN) architecture capable of simultaneously fooling both the SV and PAD modules. Trained on bona fide and spoofed samples, the model generated high-quality synthetic utterances that significantly degraded system performance. Evaluations on ASVspoof 2019 showed that SV accuracy dropped below 20% while PAD effectiveness was severely compromised, illustrating the threat posed by jointly optimized adversarial audio.
VII-B Black-Box Attacks
To overcome the limitations of white-box assumptions, several studies have explored black-box strategies that do not require gradient access.
Xuan Hai et al. [101] presented SiFDetectCracker, a query-based black-box attack that perturbs non-speech segments (e.g., silence, background noise) to exploit the reliance of some PAD systems on environmental cues. The attack achieved a success rate of over 80% without altering speaker-identifying vocal components, ensuring the audio was still accepted by the ASV system. This highlights that adversarial success does not require manipulating the entire signal—targeting peripheral acoustic features can be sufficient.
A more physically robust approach was introduced by Kassis et al. [102], who proposed a lightweight signal processing attack that successfully deceived both ASV and CM systems during over-the-air playback. By cloning prosodic and spectral features from high-quality speaker recordings, the attack bypassed multiple detection systems. However, it was less effective against advanced models, such as Wav2Vec2-based PADs. The attack also assumes the attacker can inject audio via a rooted phone, a prerequisite often blocked by the security measures of high-stakes applications.
Another category of black-box attacks involves creating universal adversarial filters. Panariello et al. [103] introduced Malafide, a fixed Finite Impulse Response (FIR) filter learned via adversarial training. This universal filter can be convolved with any spoofed utterance to evade detection, removing the need to craft input-specific perturbations. This method proved effective against high-performing deep PADs and is practical for real-time applications.
This approach was extended in Malacopula [104], which created a nonlinear filter targeting the ASV component itself by enhancing spoofed audio to better match the victim’s voiceprint. These methods demonstrate a clear trend toward physically realizable adversarial strategies that can survive digital-to-analog conversion and transmission through real-world channels.
VIII Evaluation
TABLE VI: Comparison of Voice Authentication and Spoofing Datasets.
Dataset Purpose # Spkrs Content Type Spoofing Attacks
YOHO [70] SV 138 Fixed phrases, mic-recorded in lab
VoxCeleb1/2 [106, 107, 58] SV 1,251 / 6,112 Celebrity interviews (YouTube)
AISHELL1/2 [108, 109] Mandarin SV 400+ Mandarin, read speech
TIMIT [110] Phonetics / SID 630 Phonetic, scripted, studio speech
NTIMIT [111] SR (Noisy Env.) 630 Phonetic speech, telephone-quality
LibriSpeech [112] ASR / SID $\sim$2,400 Audiobooks (clean/noisy)
Common Voice [113] Multilingual SR >>90,000 Crowdsourced, multilingual speech
VCTK [114] VC / SID 109 Read speech, accented, studio-recorded
SVCC [115] VC benchmark Varies Natural and converted speech VC
ASVspoof 2015 [116] Anti-spoofing 106 Logical Synth., Replay
ASVspoof 2017 [117] Replay detection 10 Replay across acoustic setups Replay
ASVspoof 2019 [105] Advanced spoofing 121 Logical (VC/TTS), Physical (Replay) VC, TTS, Replay
ASVspoof 2021 [118] Deepfake detection 208 Neural (VC/TTS), replay, noisy cond. Neural VC, TTS, Replay
FoR [119] Deepfake detection 110 Real vs. TTS clips TTS
BackdoorVoice [120] Data poisoning 150 Clean vs. triggered utterances Backdoor
VIII-A Popular Datasets
Evaluating VAS and studying their vulnerabilities requires access to diverse and well-annotated datasets. Over the past decade, several benchmark datasets have been developed to support research in SV, spoofing detection, and adversarial robustness. An overview of dataset distribution per field is illustrated in Figure 5, offering a visual breakdown of the number of datasets categorized by their primary research purpose. Additionally, A detailed comparison of these datasets is provided in Table.  VI.
Figure 5: Distribution of voice-related datasets across different research fields such as Speaker Verification, Spoofing Detection, Voice Conversion, Speech Recognition, and Data Poisoning.
Among the most widely used for SV are VoxCeleb1 [106] and VoxCeleb2 [107], which consist of real-world audio samples collected from celebrity interviews on YouTube. These datasets are particularly valuable for training and evaluating deep learning models in unconstrained, noisy environments. LibriSpeech [112], TIMIT [110], and NTIMIT [111] offer cleaner and more controlled conditions, with NTIMIT providing a telephone-quality version of TIMIT that introduces channel variability while preserving phonetic alignment. This makes them suitable for baseline models, phonetic analysis, and exploring channel robustness. Mozilla’s Common Voice [113] and the AISHELL datasets [108, 109] expand the focus to multilingual and Mandarin speech, respectively.
In the context of spoofing and synthetic voice detection, the ASVspoof challenges  [116, 117, 105, 118] have provided standardized protocols and data for evaluating systems under various types of attacks, including replay, TTS, and VC. Recently, datasets such as Fake or Real (FoR) [119] and BackdoorVoice [120] are specifically designed for deepfake detection and data poisoning attack evaluation, respectively. The Voice Conversion Challenge (SVCC) datasets [115] also contain a diverse set of converted and genuine voice samples to benchmark VC systems and anti-spoofing models.
VIII-B Evaluation Metrics
The standard metrics for evaluating SV systems include:
  • Equal Error Rate (EER): The rate at which the false acceptance rate (FAR) equals the false rejection rate (FRR). Lower EER indicates better system accuracy.
  • Detection Error Tradeoff (DET) Curve: A plot showing the tradeoff between FAR and FRR over different thresholds, used to visualize system performance.
  • Attack Success Rate (ASR): The proportion of adversarial attempts that successfully fool the system. A higher ASR indicates the system is more vulnerable.
  • Signal-to-Noise Ratio (SNR): Measures the strength of the original signal relative to adversarial noise. Higher SNR indicates less perceptible attacks.
    $\text{SNR (dB)}=10\log_{10}\left(\frac{P_{\text{signal}}}{P_{\text{noise}}}\right)$
  • t-DCF (tandem Detection Cost Function): A primary metric for evaluating automatic speaker verification systems in combination with countermeasure systems. Lower t-DCF values indicate better joint system performance.
    $\text{t-DCF}=C_{\text{miss}}^{\text{ASV}}\cdot\pi_{\text{tar}}\cdot P_{\text{miss}}^{\text{ASV}}+C_{\text{fa}}^{\text{ASV}}\cdot\pi_{\text{non}}\cdot P_{\text{fa}}^{\text{ASV}}+C_{\text{fa}}^{\text{CM}}\cdot\pi_{\text{spoof}}\cdot P_{\text{fa}}^{\text{CM}}+C_{\text{miss}}^{\text{CM}}\cdot\pi_{\text{tar}}\cdot P_{\text{miss}}^{\text{CM}}$
VIII-C Competitions
Table VII provides a comparative overview of the major anti-spoofing competitions, summarizing their key characteristics, attack types, and contributions to the field.
TABLE VII: Summary of Major Anti-Spoofing Competitions
Competition Year Attacks Focus Key Contribution
ASVspoof 2015 [116] 2015 VC, SS Standalone ASV spoofing Introduced standard protocols for synthetic spoofing detection
ASVspoof 2017 [117] 2017 Replay Physical replay attacks Real-world replay attacks recorded in diverse acoustic conditions
ASVspoof 2019 [105] 2019 VC, SS, Replay Logical and physical access Used neural vocoders for synthesis; first division into LA and PA tasks
DFDC Audio Track [121] 2020 Deepfake (multi-modal) Multimodal fake detection Included manipulated audio in conjunction with manipulated video
ASVspoof 2021 [118] 2021 LA, DF, VC Generalized anti-spoofing Focused on generalization across spoofing types and unseen attacks
Microsoft ADD [122] 2022 TTS, VC, Deepfake Audio deepfake detection Realistic multi-accent dataset targeting fake audio from TTS/VC
VIII-C1 Speaker Verification Competitions
Several benchmark competitions have significantly improved the field of speaker verification. The NIST Speaker Recognition Evaluations (SRE) [123] have played a critical role since 1996 in shaping speaker recognition research. These evaluations focus on assessing systems under a wide range of real-world conditions, including language variability, channel effects, and environmental noise.
More recently, the VOiCES from a Distance Challenge [124] has targeted speaker recognition in far-field and acoustically challenging environments, using data captured in realistic room settings. Such challenges push the boundaries of SRS in practical, low-SNR conditions.
VIII-C2 Anti-spoofing Competitions
With the rise of spoofing threats, dedicated anti-spoofing competitions have emerged. The ASVspoof Challenge series [116, 117, 105, 118] is the most prominent, providing standardized protocols and datasets to evaluate spoof detection systems. Each edition of ASVspoof introduces new attack types, evolving from replay and voice conversion to advanced synthetic speech generated by neural vocoders and TTS systems.
Other initiatives include the Microsoft ADD Challenge [122], which broadens the scope to deepfake audio detection, and the Deepfake Detection Challenge (DFDC) [121], which, although primarily visual, includes audio-based manipulations in its multi-modal dataset.
IX Comparison and Limitations
This section provides a comparative analysis of the attack methods discussed throughout the previous sections. Table VIII summarizes the common limitations identified across all reviewed attack categories, highlighting shared vulnerabilities and constraints that hinder their practical deployment in real-world voice authentication systems.
IX-A Data Poisoning
Untargeted data poisoning methods like Our-UE [40], WaveFuzz [41], and SyntheticPop [42] rely on clean-label manipulation to degrade model performance stealthily. While effective, they typically require high poisoning ratios ($\geq$10%) and suffer from limited scalability, generalizability, and robustness to preprocessing.
Targeted attacks, such as those by Li et al. [44] and Mohammadi et al. [45], aim to induce specific misclassifications with minimal global impact. Though successful under ideal conditions, they often rely on unrealistic assumptions like full control over training data or victim-labeled injections, limiting real-world applicability.
Backdoor attacks exploit hidden triggers to activate malicious behavior without degrading clean performance. Advanced methods like EmoBack [51], PBSM [49], and FakeVC [50] achieve high success with minimal poisoning, but they face challenges in stealth, trigger reliability, and transferability across architectures or environments.
Across all categories, major limitations persist: reliance on white-box assumptions, sensitivity to audio preprocessing, poor scalability, and environmental fragility. While defenses (e.g., pruning, robust training) show promise, they often reduce accuracy or add complexity. Future research should focus on attacks that are adaptive, low-resource, and resilient in real-world, black-box conditions.
IX-B Adversarial Attacks
Among the studies reviewed in Section V, adversarial attacks on SRSs largely focus on white-box and black-box scenarios. White-box methods leverage model gradients, while black-box methods aim for query-efficient, imperceptible attacks. Despite innovations such as psychoacoustic modeling and ensemble-based surrogates, practical limitations remain due to reliance on score access, high query budgets, or constrained physical-world setups.
Physical-world attacks and evaluations under realistic conditions, such as PhyTalker[68] and AS2T[67], are still uncommon and often assume semi-controlled environments. Most studies overlook countermeasure-resilient or fully over-the-air attack settings, which limits their real-world use. Additionally, generalization across different languages, time changes, and various conditions is not adequately covered. Most evaluations focus on specific datasets and assume that acoustic conditions remain unchanged.
On the defense side, progress lags significantly. Aside from a few survey mentions, proactive or real-time defenses are underexplored, and existing countermeasures are largely reactive and lack standard benchmarking. This highlights a growing asymmetry between offensive advancements and defensive maturity, signaling a need for more robust, deployment-ready protections.
TABLE VIII: Limitations Comparison for Data Poisoning, Adversarial, Deepfake, and Adversarial Spoofing Attacks
Limitation Category Data Poisoning Attacks Adversarial Attacks Deepfake Attacks Adversarial Spoofing Attacks
Requires model or training access ✓ High $\triangle$ Partial ✗ None $\triangle$ Partial
Breaks under preprocessing ✓ Yes ✓ Yes ✓ Yes ✓ Yes
Low transferability ✓ Present ✓ Present ✓ Present ✓ Present
Needs high resource/querying $\triangle$ Sometimes ✓ Yes ✓ Often ✓ Often
Limited real-world robustness ✓ Not robust ✓ Not robust ✓ Not robust ✓ Not robust
Multilingual/generalization gaps ✓ Present ✓ Present ✓ Present ✓ Present
Evaluated in ideal conditions ✓ Mostly ✓ Mostly ✓ Mostly ✓ Mostly
Scalability challenges ✓ Significant ✓ Significant ✓ Significant ✓ Significant
IX-C Deepfake Attacks
As outlined in Section VI, deepfake-based attacks present an increasingly serious threat to SRSs, leveraging generative models for voice imitation or partial utterance replacement. Unlike traditional adversarial attacks that rely on optimization or model access, deepfake methods often succeed in fully black-box, zero-shot settings—making them highly practical and scalable. Early work using voice morphing and more recent partial spoofing (PF) techniques confirm the vulnerability of both machine and human verification systems, even with minimal attacker knowledge.
However, these attacks face limitations in real-world deployment. Most are evaluated in ideal, noise-free settings and rarely consider multilingual, low-quality, or dynamic acoustic conditions. Furthermore, generative artifacts and timing inconsistencies remain detectable in some models, and defense-aware strategies are largely unexplored. Compared to adversarial audio, where countermeasure bypass is an active research area, deepfake studies often neglect robust CM evaluations.
Overall, deepfake and adversarial approaches expose complementary weaknesses in SRSs—optimization-based attacks exploit model internals, while generative synthesis bypasses through realism. This calls for unified defense mechanisms capable of detecting both perturbation- and synthesis-based threats amid evolving generative audio capabilities.
IX-D Adversarial Spoofing Attacks
As discussed in Section VII, most reviewed defenses rely on passive, model-agnostic strategies, yet they fall short against modern, adaptive attack techniques. Recent adversarial methods, discussed in Section V, exploit subtle, non-obvious audio features or leverage universal perturbations that evade traditional PAD systems—even under physical conditions. Notably, none of the defenses evaluated address real-time constraints or defense latency, leaving systems vulnerable to fast, transferable attacks.
White-box attacks demonstrate the fragility of PADs under full access but face scalability and perceptual quality limitations. In contrast, black-box and signal-level attacks offer more realistic threat models but often depend on query access or suffer from environmental sensitivity, limiting their deployment potential. While GAN-based approaches can deceive end-to-end systems with high perceptual quality, they require intensive training and remain sensitive to architectural variance.
Across the board, key gaps persist: limited evaluation in multilingual, cross-device, or live settings, and a lack of defense solutions that are both attack- and model-agnostic. These gaps underline the urgent need for adaptive, proactive PAD systems capable of low-latency, robust detection in dynamic real-world scenarios.
X Future Work
Despite significant progress in voice-based biometric systems, emerging threats continue to challenge the robustness, generalizability, and real-world applicability of current defenses. We structure this section around four major threat types: data poisoning, adversarial attacks, deepfake attacks, and adversarial spoofing attacks.
Several challenges, such as limited transferability, unrealistic evaluations, and deployment inefficiency, are common across these threats. These issues are discussed in each relevant subsection and then summarized at the end.
X-A Data Poisoning Attacks
  1. 1.
    Detection of Stealthy Poisoning: Current defenses struggle to detect ultra-subtle manipulations affecting ¡1% of training data. Future work should explore anomaly-sensitive detection using unsupervised learning or contrastive objectives.
  2. 2.
    Realistic Threat Models: Most poisoning studies assume white-box access and full control. Research should consider practical environments, such as federated learning, data marketplaces, or API-based systems.
  3. 3.
    Transferability of Poisoned Data: The persistence of poisoning effects across architectures and datasets remains underexplored. Understanding this will improve generalization and strength.
  4. 4.
    Low-Cost and Adaptive Attacks: Existing poisoning methods are often resource-intensive. Future attacks should be lightweight and able to change in response to defenses.
  5. 5.
    Hybrid Attack Pipelines: Combining poisoning with adversarial inputs at test time can create complex pipelines. It is important to formalize and defend against these for overall system security.
X-B Adversarial Attacks
  1. 1.
    Universal and Transferable Perturbations: Adversarial attacks often fail across architectures and datasets. Future efforts should develop universal perturbations via meta-learning or training-agnostic methods.
  2. 2.
    Realistic Benchmarking: Many attacks are evaluated on static or clean inputs. Evaluations should include noise, compression, spontaneous dialogue, and differences between devices.
  3. 3.
    Robustness in the Physical World: Adversarial examples degrade under real-world conditions. Research into real-time, dynamic perturbations can improve attack effectiveness in live settings.
  4. 4.
    Query-Efficient Black-Box Attacks: Current methods are often impractical due to query limits. Decision-based or generative surrogate models with little supervision are promising alternatives.
  5. 5.
    Cross-Lingual Transferability: Adversarial performance drops sharply outside English datasets. Research in this area is still very limited, and future work should explore language-agnostic and multilingual attack techniques.
X-C Deepfake Attacks
  1. 1.
    Robust Detection in Real-World Environments: Detection models must work under telephony, reverberation, crosstalk, and overlapping speech conditions, often ignored in current benchmarks.
  2. 2.
    Cross-Language and Accent Generalization: Most models are trained on English. We need multilingual datasets and features that work across different accents for global use.
  3. 3.
    Artifact-Based Detection: Deepfakes show subtle inconsistencies in their waveforms. Future detection should use frame-level alignment, waveform modeling, and detailed spectral cues.
  4. 4.
    Efficient and Real-Time Inference: Models must be deployable on edge devices and operate in real-time. Techniques such as quantization and early-exit models are needed.
  5. 5.
    Adaptive Deepfakes and Defense Evasion: Future evaluations should simulate adversarial deepfakes trained to evade known detectors, to benchmark true model robustness.
X-D Adversarial Spoofing Attacks
  1. 1.
    Threat Model Standardization: Evaluation assumptions are inconsistent. Clear definitions for white-box, gray-box, and black-box setups will improve reproducibility.
  2. 2.
    Adversarial Training of CMs: Most CMs are not robust to adversarial audio. Training with perturbed examples is necessary to strengthen their resilience.
  3. 3.
    Generalizable Detection Strategies: CMs often overfit to specific spoofing types. Broader methods, like signal fingerprinting or multi-modal biometric fusion, can improve generalization.
  4. 4.
    Active Defense Mechanisms: Passive defenses can be bypassed. Active strategies such as randomized prompts or timing jitter can disrupt attack optimization.
  5. 5.
    Low-Latency and Edge-Ready Defenses: Deployment requires real-time capabilities. Early-exit scoring, frame-level anomaly detection, and lightweight ensembles are viable paths.
Similar Challenges
Many of the challenges mentioned above are common across different threat categories. We list the most important ones below:
  1. 1.
    Transferability: Attacks and defenses often lack generalization across models, datasets, and languages. Developing universal perturbation and detection strategies remains a critical research frontier.
  2. 2.
    Realism in Evaluation: Benchmarks frequently ignore real-world artifacts such as compression, ambient noise, spontaneous dialogue, or cross-device variability. Future evaluation protocols must simulate realistic deployment settings.
  3. 3.
    Efficiency and Deployment Constraints: Both attack and defense mechanisms must be optimized for real-time, low-power, and bandwidth-constrained environments, especially for mobile and edge devices.
  4. 4.
    Standardization of Threat Models: Inconsistent definitions of white-box, gray-box, and black-box scenarios hinder reproducibility and fair benchmarking. Future work should adopt shared assumptions and standardized evaluation setups.
Addressing these cross-cutting challenges is crucial to secure voice authentication systems and ensure they can withstand changing threats.
XI Conclusion
In this survey, we provided a comprehensive review of the threats targeting voice authentication systems. We categorized and analyzed the main types of attacks, including data poisoning, adversarial, deepfakes, and adversarial spoofing attacks. For each threat, we summarized the key techniques, identified commonly used datasets, and examined the main taxonomies used in the field. We conducted detailed comparisons and highlighted limitations across the different attack strategies. Furthermore, we organized existing research based on widely adopted taxonomies and identified key open challenges. Finally, we outlined future research directions to guide and inspire continued advancements in securing voice authentication systems.
Acknowledgments
This research is supported by the Air Force Office of Scientific Research under award number FA2386-23-1-4003.